OpenStudy (anonymous):

Given that a person married in 2010 was less than twenty years old, what is the probability that the person was male? Given that a person married in 2010 was female, what is the probability that the person was between the ages of 35 and 44? If a person that was married in 2010 was randomly chosen, what is the probability that the the person is male and between the ages of 25 and 34? Less than 20 20-24 25-34 35-44 45 + Total Male 913 9847 23923 9707 9792 54182 Female 2110 13500 22581 8386 7605 54182 Totals 3023 23347 46504 18093 17397 108364

OpenStudy (anonymous):

a). P(person) = P(male<20)/ P(totals total) = 913/108364 b). P(female) = P(female 35-44)/P(totals total) = 8386/108364

OpenStudy (anonymous):

you are awesome thanks now I can actually do the homework and know what I am doing. Thank you so so so much